# Can Large Language Models Automatically Score Proficiency of Written Essays?

### v1

@@ 1.   Introduction
Recently, Large Language Models (LLMs) were introduced as cutting-edge transformer-based models demonstrating impressive capabilities on various tasks (Meoni et al., 2023; Kocmi and Federmann, 2023). In this paper, we check if the language proficiency retained by LLMs (represented by ChatGPT [^1] https://chat.openai.com/chat and Llama 2222https://ai.meta.com/llama/ in our study) is useful in scoring essays. We test the capabilities of LLMs in this task across two levels, holistically and per individual writing trait. We also compare its performance to current state-of-the-art (SOTA) models. Furthermore, we analyze its performance and check its consistency in scoring the same set of essays.
@@
Our developed prompts led to notable enhancements in the performance of LLMs, reaching a peak QWK score of 0.606 and 0.562 by ChatGPT and Llama, respectively. Although those scores lag behind SOTA models with respect to scoring prediction, LLMs have an advantage in terms of generated feedback. For all the experimented essays, they provided feedback to enhance their quality.
@@
  • We study the consistency of ChatGPT and Llama in AES.
@@ 2.   Related Work
Previous attempts to address the AES problem relied either on crafting features to represent essays, syntactically Yannakoudakis et al. (2011) or lexicallay Chen and He (2013); Phandi et al. (2015), or adopting neural networks Xie et al. (2022). As a matter of fact, SOTA performance was achieved using a set of Multi-layer perceptrons that simultaneously conducts regression and ranking optimization Xie et al. (2022).
Following the advancement of neural networks, LLMs were utilized in various applications. Consequently, a handful of researchers explored fine-tuning LLMs like BERT Devlin et al. (2018) for the AES task, which yielded a remarkable performance Wang et al. (2022); Yang et al. (2020).
ChatGPT and Llama [^5] Hereafter we will refer to Llama 2 as Llama. were built using reinforcement learning from human feedback to capture contextual meanings in a much deeper way, leading to a better language generation. Those models can be utilized for multiple tasks, e.g., code generation and text paraphrasing. To the best of our knowledge, the only study that employed a GPT-based model for AES was conducted by Mizumoto and Eguchi (2023). Utilizing OpenAI’s text-davinci-003 model for text completion, they automatically scored the essays of the ETS Corpus of Non-Native Written English (TOEFL11 dataset) and employed the generated scores as features of a Bayesian regression model. They also combined these scores with a range of linguistic features, exploring their influence on performance. That work, however, did not position the results with previous work. Moreover, it was limited to holistic scoring with no exploration of per-trait performance, in addition to experimenting with one short prompt without fully exploiting the potential of prompt engineering to enhance the prompt’s structure and clarity.
In this paper, we aim to assess the capability of using LLMs in AES over two levels: holistically and per trait. Moreover, different prompts are crafted to study the effect of prompt engineering on the performance of LLMs.
@@ 3.   Methodology of Study
This section presents the outline of the adopted model and dataset, the design of the experimented prompts, and the experimental design.
@@ 3.1.   Dataset and Models
Task Essay Type
Avg length
Score Range # Essays
1 Persuasive
350
2-12 1783
2
350
1-6 1800
3 Source-Dependent
150
0-3 1726
4
150
0-3 1772
5
150
0-4 1805
6
150
0-4 1800
7 Narrative
250
0-30 1569
8
650
0-60 723
Table 1: Properties of the different tasks in ASAP dataset.
We have chosen two LLMs models for our experiments, namely, ChatGPT and Llama. ChatGPT has demonstrated remarkable capabilities on various tasks and gained significant attention across diverse domains. In our experiments, we utilized the ChatGPT API, namely gpt-3.5-turbo-0301 (a Snapshot from March 1st, 2023). [^6] https://platform.openai.com/docs/models This model is optimized for dialogue, and its performance is on par with Instruct Davinci. Nevertheless, ChatGPT API is not open source, which imposes scalability limitations due to its associated cost. [^7] The total cost of the experiments with ChatGPT in this work amounted to approximately \$120. Alternatively, Llama is an open-source LLM developed by Meta. It outperforms other open-source language models (e.g., Falcon and MPT) on different benchmarks. In this work, we used Llama-2-13b-chat-hf [^8] https://huggingface.co/meta-llama/Llama-2-13b-chat-hf. For both models, the temperature is set to 0 to guarantee reproducible results.
@@ 3.2.   Prompt Design
Ensuring the clarity of the prompt given to the LLMs is vital. Therefore, we follow some prompt-engineering tactics advised by OpenAI [^10] https://platform.openai.com/docs/guides/gpt-best-practices/strategy-write-clear-instructions to enhance the prompt comprehension. First, proper delimiters are used to distinguish the inputs (essay, rubrics, and one-shot examples) from the instruction text, preventing prompt injection issues. Second, the prompt is designed as a series of sequential instructions, pushing the models to complete the steps in order. Third, we instruct the models to preprocess the input text (as shown in Figure 4), as certain placeholders are present within ASAP essays. Finally, we ask for a JSON output structure specifying the desired outputs. Overall, we designed four prompt variants that meet these rules as follows:
  • Prompt A (shown in Figures 1 and 2) is a simple prompt asking to evaluate the input essay without defining the rubrics. It anchors our expectations on the performance of ChatGPT and Llama without any extra information.
    Your job is to evaluate the provided essay on a scale from 1 to 6. {preprocessing_instruction} Here is a summary of the required steps: 1. Evaluate the essay (after suitable replacement of the anonymization tokens wherever needed) on a scale from 1 to 6. 2. Format your response in JSON as follows: { “Total_score”: your evaluation score, “Commentary”: your feedback on the essay } Here is the prompt delimited by <>: <Prompt text> And here is the essay, which you need to evaluate, delimited by triple backticks: “‘input_essay“‘
    Figure 1: Prompt A, a simple prompt asking ChatGPT to score an essay according to the score range.
@@
The prompts for ChatGPT and Llama contain the same instructions but in different formats. We tried multiple arrangements of the instructions inside the prompt to see the best instruction order for each LLM. After several trials with the order of the instructions inside the prompts, we noticed two main differences. First, ChatGPT benefits from adding a concise list of required steps, whereas Llama did not exhibit similar enhancements from these steps. Second, ChatGPT provided the required JSON format, which was one of the required steps, regardless of the position of this requirement inside the prompt. On the other hand, with Llama, the required format is only generated when explicitly specified at the very end of the prompt. Another difference in the prompts between ChatGPT and Llama is that in Llama’s prompts, we did not ask for feedback. We did this to keep the responses concise because Llama often offers extra feedback even when it is not asked to provide one. Figures 1 and 2 show the structure of prompt A for ChatGPT and Llama. Prompts B and C have similar formats with the addition of the rubric and the one-shot example.
Figure 4: The preprocessing instruction
@@
Considering ChatGPT’s paid API, testing all tasks on 4 prompts is quite costly. Given that ASAP tasks are divided into three categories, we performed the experiments in two phases for both ChatGPT and Llama. In the first phase, we experiment with one task from each category on all of our prompts to choose the best prompt for each category. In the second phase, we apply that prompt to the remaining tasks in that category. From each category, we opt for the task with the shortest overall length (rubric guidelines, essay resource, and task question). So, we selected task 1 (persuasive), task 5 (source-dependent), and task 7 (narrative) for the first phase.
@@ 4.   Experimental Results and Discussion
In this Section, we discuss the performance of ChatGPT and Llama in the AES task. Moreover, we provide a thorough examination of the differences between ChatGPT and Llama when it comes to judging essay proficiency. Recall that in this work, we aim to answer the following research questions:
@@
Our baselines are SOTA on both the holistic and trait levels. As such, we contrast our LLM models with Xie et al. (2022); Jiang et al. (2023) from a holistic perspective and against Mathias and Bhattacharyya (2020) over trait level. For the holistic level, we chose two baselines, which achieved SOTA results for in-prompt AES, and cross-prompt AES, Xie et al. (2022) and Jiang et al. (2023) respectively. These baselines show us the highest agreement neural models reached when they were trained on a specific task and tested on it, and when they were trained on different tasks and tested on an unseen task.
4.1.   Effect of Prompt Engineering (RQ1 & RQ2)
@@ A B C D
Table 2 presents the results of the first phase of experiments, which reveal several observations.
@@
From this analysis, Llama demonstrated inconsistent performance in AES, which suggests that it is highly sensitive to prompt selection. ChatGPT showed higher agreement across the different prompts and tasks, indicating that ChatGPT exhibits consistent performance in AES. It is worth noting that in the ASAP dataset, the average agreement score between the two raters is 0.75  [^12] For rater1_domain1 and rater2_domain1, which indicates the complexity of the AES task. These results show the critical role of prompt engineering in enhancing or diminishing the performance of LLMs on the AES task. Moreover, choosing the right prompt is highly dependent on both the LLM and the nature of the task.
@@
Table 4 compares the performance of ChatGPT and Llama per ASAP task with SOTA models (as reported in their respective studies) on the holistic level. First, we notice that the average performance of the two LLMs across tasks is quite comparable, with a slight advantage for ChatGPT. The results also show that ChatGPT does a reasonably good job in source-dependent tasks 4-6, where the performance is significantly better than in the other (source-independent) tasks, while Llama showed a better performance only in Task 2 (which is persuasive). Second, it is quite apparent that the exhibited performance of both LLMs is nowhere near the SOTA models, with average QWK of 0.817, 0.794, 0.313, and 0.297 for Xie et al. (2022); Yang et al. (2020), ChatGPT, and Llama respectively. While those SOTA models are task-specific and trained on the ASAP dataset directly, higher performance was expected from the LLMs, given the vast amount of data they were trained on and their ability to generate good-quality text. It is clear that although LLMs were able to understand the task and give a score, it was not able to capture the input essay as a whole and score it accordingly.
@@
On the level of the trait, Table 5 draws a comparison with SOTA models per trait for tasks 7 and 8, the only tasks that have traits in ASAP. The results deliver the same message of the holistic level; ChatGPT and Llama are very far behind.
4.3.   Analysing the Provided Feedback
@@
Essay
ChatGPT
Llama
@@
The response is completely irrelevant and incorrect. It does not address the prompt and does not demonstrate any understanding of the text.
The response shows evidence of a minimal understanding of the text. The essay does not address the prompt or the features of the setting that affect the cyclist. Instead, the essay focuses on the writer’s personal feelings and experiences, which are not related to the prompt. The essay also contains some inaccuracies, such as the statement that the cyclist was “cut hurt” and that the features of the setting did not affect the cyclist. Overall, the response lacks a clear understanding of the text and the prompt.
@@
The response demonstrates a partial understanding of the text. The essay discusses how the features of the setting affect the cyclist, specifically how the old directions and abandoned towns with no good water could have killed him. However, the response could have provided more examples from the essay to support the conclusion. Additionally, the response could have extended the understanding beyond the literal by discussing how the cyclist’s experience reflects the fluidity of the world and the importance of being prepared for nature’s challenges.
The response demonstrates a partial understanding of the text, although may not develop all parts equally. The essay provides some examples from the text to support the conclusion, but the analysis is not comprehensive and does not fully connect the support to a conclusion or assertion made about the text. The response correctly identifies the features of the setting that affect the cyclist, such as the desert environment, the lack of water, and the abandoned towns. However, the analysis could be more detailed and nuanced. For example, the response could discuss how the setting affects the cyclist’s mood and confidence, or how the abandoned towns and lack of water create a sense of desolation and hopelessness. Additionally, the response could provide more specific examples from the text to support the analysis. Overall, the response demonstrates some understanding of the text, but could be more comprehensive and nuanced in its analysis.

### v2

@@ 1.   Introduction
Recently, Large Language Models (LLMs) were introduced as cutting-edge transformer-based models demonstrating impressive capabilities on various tasks (Meoni et al., 2023; Kocmi and Federmann, 2023). In this paper, we check if the language proficiency retained by LLMs (represented by ChatGPT [^1] https://chat.openai.com/chat and Llama 2222https://ai.meta.com/llama/ in our study) is useful in scoring essays. We test the capabilities of LLMs in this task across two levels, holistically and per individual writing trait. We also compare its performance to current state-of-the-art (SOTA) models. Furthermore, we analyze its performance and check its consistency in scoring the same essays.
@@
Our developed prompts led to notable enhancements in the performance of LLMs, reaching a peak Quadratic Weighted Kappa Cohen (1968) score of 0.606 and 0.562 by ChatGPT and Llama 2, respectively. Although those scores lag behind SOTA models with respect to scoring prediction, LLMs have an advantage in terms of generated feedback. For all the experimented essays, they provided feedback to enhance their quality.
@@
  • We study the consistency of ChatGPT and Llama 2 in AES.
@@ 2.   Related Work
Previous attempts to address the AES problem relied either on crafting features to represent essays, syntactically Yannakoudakis et al. (2011) or lexicallay Chen and He (2013); Phandi et al. (2015), or adopting neural networks Xie et al. (2022). As a matter of fact, SOTA performance was achieved using a set of Multi-layer perceptrons that simultaneously conduct regression and ranking optimization Xie et al. (2022).
Following the advancement of neural networks, LLMs were utilized in various applications. Consequently, a handful of researchers explored fine-tuning LLMs, e.g., BERT Devlin et al. (2018), for the AES task, which yielded a remarkable performance Wang et al. (2022); Yang et al. (2020).
ChatGPT and Llama [^5] Hereafter we will refer to Llama 2 as Llama. were built using reinforcement learning from human feedback to capture contextual meanings in a much deeper way, leading to a better language generation. Those models can be utilized for multiple tasks, e.g., code generation and text paraphrasing. Although work utilizing LLMs for AES is limited, the field is quickly evolving. Existing work solely focus on proprietary models. For example, Mizumoto and Eguchi (2023) utilized OpenAI’s text-davinci-003 model for text completion to automatically score the essays of the TOEFL11 dataset and employed the generated scores as features for a Bayesian regression model. naismith-etal-2023-automated tested GPT-4 achiam2023gpt on assessing the discourse coherence of essays, yielding promising results. Similarly, yancey-etal-2023-rating focused on GPT-3.5 and GPT-4 with Common European Framework of Reference for Languages (CEFR) short essays dataset. Both models fell short behind a simple XGBoost baseline.
Although prior research highlighted the underperformance of LLMs, some gaps persist. Firstly, existing studies focus solely on OpenAI’s paid models. Secondly, none of them have tested these models on the ASAP dataset, an essential benchmark in AES for comparing with SOTA models.
In this paper, we aim to assess the capability of using LLMs in AES over two levels, holistically and per trait. Moreover, different prompts are crafted to study the effect of prompt engineering on the performance of LLMs.
@@ 3.   Methodology of Study
This section presents the outline of the adopted models and dataset, the design of the experimented prompts, and the experimental design.
@@ 3.1.   Dataset and Models
We have chosen two LLM models for our experiments, namely, ChatGPT and Llama. ChatGPT has demonstrated remarkable capabilities on various tasks and gained significant attention across diverse domains. In our experiments, we utilized the ChatGPT API, namely gpt-3.5-turbo-0301 (a Snapshot from March 1st, 2023). [^6] https://platform.openai.com/docs/models This model is optimized for dialogue, and its performance is on par with Instruct Davinci. Nevertheless, ChatGPT API is not open source, which imposes scalability limitations due to its associated cost. [^7] The total cost of the experiments with ChatGPT in this work amounted to approximately \$120. Alternatively, Llama is an open-source LLM developed by Meta. It outperforms other open-source language models (e.g., Falcon and MPT) on different benchmarks Touvron et al. (2023). In this work, we used Llama-2-13b-chat-hf. [^8] https://huggingface.co/meta-llama/Llama-2-13b-chat-hf For both models, the temperature is set to 0 to guarantee reproducible results.
{tblr}
colspec = cp0.9cmcp0.8cmc, cells = c, cell22 = r=2, cell42 = r=4, cell82 = r=2, hline1-2,4,8,10 = -, Task & Essay Type Avg. length Score Range # Essays
1 Persuasive 350 2-12 1783
2 350 1-6 1800
3 Source-Dependent 150 0-3 1726
4 150 0-3 1772
5 150 0-4 1805
6 150 0-4 1800
7 Narrative 250 0-30 1569
8 650 0-60 723
Table 1: Properties of the different tasks in the ASAP dataset.
Note that the essay might contain some anonymization tokens that are placeholders added to anonymize sensitive information and ensure privacy. They replace personal names, locations, numbers, times, dates, organizations, and other specific information that could identify individuals or institutions. The anonymization tokens are: @PERSON#, @LOCATION#, @NUM#, @TIME, @MONTH, @DATE, and @CAPS#, where # refers to a number.
Figure 1: The preprocessing instructions.
@@ 3.2.   Prompt Design
Ensuring the clarity of the prompt given to the LLMs is vital. Therefore, we follow some prompt-engineering tactics advised by OpenAI [^10] https://platform.openai.com/docs/guides/gpt-best-practices/strategy-write-clear-instructions to enhance the prompt comprehension. First, proper delimiters are used to distinguish the inputs (essay, rubrics, and one-shot examples) from the instruction text, preventing prompt injection issues. Second, the prompt is designed as a series of sequential instructions, pushing the models to complete the steps in order. Third, we instruct the models to preprocess the input text (as shown in Figure 1), as certain placeholders are present within the ASAP essays. Finally, we ask for a JSON output structure specifying the desired outputs. Overall, we designed four prompt variants that meet these guidelines as follows:
  • Prompt A (shown in Figures 2 and 3) is a simple prompt asking to evaluate the input essay without defining the rubrics. It anchors our expectations on the performance of ChatGPT and Llama without any extra information. Your job is to evaluate the provided essay on a scale from 1 to 6. {preprocessing_instruction} Here is a summary of the required steps: 1. Evaluate the essay (after suitable replacement of the anonymization tokens wherever needed) on a scale from 1 to 6. 2. Format your response in JSON as follows: { “Total_score”: your evaluation score, “Commentary”: your feedback on the essay } Here is the prompt delimited by <>: <Prompt text> And here is the essay, which you need to evaluate, delimited by triple backticks: “‘input_essay“‘ Figure 2: Prompt A, a simple prompt asking ChatGPT to score an essay according to the score range.
@@
The prompts for ChatGPT and Llama contain the same instructions but in different formats. We tried multiple arrangements of the instructions inside the prompt to see the best instruction order for each LLM. After several trials with the order of the instructions inside the prompts, we noticed two main differences. First, ChatGPT benefits from adding a concise list of required steps, whereas Llama did not exhibit similar enhancements from these steps. Second, ChatGPT provided the required JSON format, which was one of the required steps, regardless of the position of this requirement inside the prompt. On the other hand, with Llama, the required format is only generated when explicitly specify it at the very end of the prompt. Another difference in the prompts between ChatGPT and Llama is that in Llama’s prompts, we did not ask for feedback. We did this to keep the responses concise because Llama often offers extra feedback even when it is not asked to provide one. Figures 2 and 3 show the structure of prompt A for ChatGPT and Llama. Prompts B and C have similar formats with the addition of the rubric and the one-shot example.
@@
Considering ChatGPT’s paid API, testing 4 prompts on all tasks is quite costly. Given that ASAP tasks are divided into three categories, we performed the experiments in two phases for both ChatGPT and Llama. In the first phase, we experiment with one task from each category on all of our prompts to choose the best prompt for each category. In the second phase, we apply that prompt to the remaining tasks in that category. From each category, we opt for the task with the shortest overall length (rubric guidelines, essay resource, and task question). So, we selected task 1 (persuasive), task 5 (source-dependent), and task 7 (narrative) for the first phase.
@@ 4.   Experimental Results and Discussion
In this Section, we discuss the performance of ChatGPT and Llama for AES. Moreover, we provide a thorough examination of the differences between ChatGPT and Llama when it comes to judging essay proficiency and their consistency. Recall that in this work, we aim to answer the following research questions:
@@
Our baselines are SOTA on both the holistic and trait levels. As such, we contrast our LLM models with Xie et al. (2022) and Jiang et al. (2023), which achieved SOTA performance for in-prompt AES and cross-prompt AES respectively, from the holistic perspective. We also compare our LLM models against Mathias and Bhattacharyya (2020) over the traits level.
4.1.   Effect of Prompt Engineering (RQ1)
@@ A B C D
Table 2 presents the results of the first phase of our experiments, which reveal several observations.
@@ 4.2.   Consistency (RQ2)
From this analysis, Llama demonstrated inconsistent performance in AES, which suggests that it is highly sensitive to prompt selection. ChatGPT showed higher agreement across the different prompts and tasks, indicating that ChatGPT exhibits consistent performance in AES. It is worth noting that in the ASAP dataset, the average agreement score between the two human raters is 0.75  [^12] For rater1_domain1 and rater2_domain1, which indicates the complexity of the AES task. These results show the critical role of prompt engineering in enhancing or diminishing the performance of LLMs on the AES task. Moreover, choosing the right prompt is highly dependent on both the LLM and the nature of the task.
@@
Table 4 compares the performance of ChatGPT and Llama per ASAP task with SOTA models (as reported in their respective studies) on the holistic level. First, we notice that the average performance of the two LLMs across tasks is quite comparable, with a slight advantage for ChatGPT. The results also show that ChatGPT does a reasonably good job in source-dependent tasks 4-6, where the performance is significantly better than in the other (source-independent) tasks, while Llama showed a better performance only in Task 2 (which is persuasive). Second, it is quite apparent that the exhibited performance of both LLMs is nowhere near the SOTA models, with average QWK of 0.817, 0.702, 0.313, and 0.297 for Xie et al. (2022), Jiang et al. (2023), ChatGPT, and Llama, respectively. While those SOTA models are task-specific and trained on the ASAP dataset directly, higher performance was expected from the LLMs, given the vast amount of data they were trained on and their ability to generate good-quality text. It is clear that although LLMs were able to understand the task and give a score, it was not able to capture the input essay as a whole and score it accordingly.
@@
On the traits level, Table 5 draws a comparison with SOTA models per trait for tasks 7 and 8, the only tasks that have traits in ASAP. The results deliver the same message of the holistic level; ChatGPT and Llama are very far behind.
$4.$   Analysing the Provided Feedback
@@
Essay ChatGPT Llama
@@
The response is completely irrelevant and incorrect. It does not address the prompt and does not demonstrate any understanding of the text. The response shows evidence of a minimal understanding of the text. The essay does not address the prompt or the features of the setting that affect the cyclist. Instead, the essay focuses on the writer’s personal feelings and experiences, which are not related to the prompt. The essay also contains some inaccuracies, such as the statement that the cyclist was "cut hurt" and that the features of the setting did not affect the cyclist. Overall, the response lacks a clear understanding of the text and the prompt.
@@
The response demonstrates a partial understanding of the text. The essay discusses how the features of the setting affect the cyclist, specifically how the old directions and abandoned towns with no good water could have killed him. However, the response could have provided more examples from the essay to support the conclusion. Additionally, the response could have extended the understanding beyond the literal by discussing how the cyclist’s experience reflects the fluidity of the world and the importance of being prepared for nature’s challenges. The response demonstrates a partial understanding of the text, although may not develop all parts equally. The essay provides some examples from the text to support the conclusion, but the analysis is not comprehensive and does not fully connect the support to a conclusion or assertion made about the text. The response correctly identifies the features of the setting that affect the cyclist, such as the desert environment, the lack of water, and the abandoned towns. However, the analysis could be more detailed and nuanced. For example, the response could discuss how the setting affects the cyclist’s mood and confidence, or how the abandoned towns and lack of water create a sense of desolation and hopelessness. Additionally, the response could provide more specific examples from the text to support the analysis. Overall, the response demonstrates some understanding of the text, but could be more comprehensive and nuanced in its analysis.